Circular No. 8068

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003ag, 2003ah, AND 2003ai Further to IAUC 8064, B. Swift, M. Ganeshalingam, and W. Li report

the LOTOSS discovery of three apparent supernovae on unfiltered KAIT images:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003ag Feb. 7.5 11^{h}26^{m}01.^{s}82 + 1^{o}59′02′′.8 16.5 12′′.5 E, 0′′.3 N
2003ah Feb. 8.2 4 43 08.54 + 0 46 00.4 17.3 7′′.1 W, 7′′.0 N
2003ai Feb. 8.5 13 00 58.68 +39 51 24.5 17.0 0′′.9 E, 7′′.7 S

Further approximate magnitude estimates: SN 2003ag in UGC 6440, 2002 Dec. 27.4 UT, [19.0; 2003 Feb. 8.3, 16.3. SN 2003ah, Jan. 29.2, [18.5; Feb. 9.2, 17.0. SN 2003ai in IC 4062, 2002 June 16.3, [19.0; 2003 Feb. 9.3, 16.9.

FLARING OBJECT IN ORION CLUSTER On behalf of a large collaborative team, K. V. Getman, E. D. Feigel-

son, and G. Garmire, Pennsylvania State University; and S. S. Murray and F. R. Harnden, Jr., Smithsonian Astrophysical Observatory, report the detection of an x-ray flare from the radio-flaring object reported on IAUC 8055 and 8060, as seen with the Chandra X-ray Observatory (+ ACIS-I detector) during an ultradeep observation of the Orion nebula during Jan. 8–22. The x-ray variability has a complex morphology: it is approximately constant around 0.04 counts/s from Jan. 8.9 to 16.3 UT, rapidly rising to 0.19 counts/s starting Jan. 16.3, falling to 0.11 counts/s by Jan. 16.7 (where it remained constant for a day), rising to 0.35 counts/s and higher over several hours starting Jan. 17.7, falling to 0.2 counts/s over one day (and remaining at that level for a half day), rising again to 0.35 counts/s and higher starting Jan. 19.7, and falling to 0.2 counts/s over one day by the end of the observation. Preliminary spectral analysis confirms the previous Chandra findings (Feigelson et al. 2002, Ap.J. 574, 258) that the source is deeply embedded in a molecular cloud with a column density of log N_{H} ∼ 22.6 cm^{−2} and with absorption-corrected luminosities of log L_{x} ∼ 31.3 and ∼ 32.0 erg/s, for the quiescent and flaring phases, respectively. As this source does not have a near-infrared excess (Muench et al. 2002, Ap.J. 573, 366), it may be an extreme example of a magnetically active weak-line T-Tau star, such as V773 Tau or DoAr 21 (cf. Feigelson and Montmerle 1985, Ap.J. 289, L19), rather than a rare radio flaring protostar (e.g., IRS 5 near R CrA; cf. Taylor and Storey 1984, MNRAS 209, 5P).

2003 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green